# Semilinear diffusion equations on infinite graphs: the dissipative and Lipschitz cases

### v1

@@ 1 Introduction
In particular, global existence and blow-up of solutions for parabolic equations has been studied by several authors in [17, 33, 39, 38, 28, 23, 41, 35].
@@
By contrast, Condition (F2) is characteristic of saturating nonlinear forcing terms, which frequently arise in biological models (see, e.g., [8]). For a detailed treatment of this case, see also [42, Chapter 31.6].
@@ 2 Preliminaries and notations
If the cardinality of the node set is finite, i.e., $|X|<\infty$, then $G$ is called a finite graph, otherwise, $G$ is called an infinite graph. The non-zero values $w(x,y)$ of the edge-weight function $w$ are called weights associated with the edge $\{x,y\}$. In this case we will write $x\sim y$ meaning that $x$ is connected to $y$. On the other hand, if $w(x,y)=0$, then we will write $x\nsim y$ meaning that $x$ and $y$ are not connected by an edge. A walk is a (possibly infinite) sequence of nodes $\{x_{i}\}_{i\geq 0}$ such that $x_{i}\sim x_{i+1}$. A path is a walk with no repeated nodes. A graph is connected if there is a finite walk connecting every pair of nodes, that is, for any pair of nodes $x,y$ there is a finite walk such that $x=x_{0}\sim x_{1}\sim\cdots\sim x_{n}=y$. Moreover, we will say that a subset $Y\subseteq X$ is connected if for every pair of nodes $x,y\in Y$ there exists a finite walk connecting $x$ and $y$ all of whose nodes are in $Y$. A subset $Y\subseteq X$ is a connected component of $X$ if $Y$ is maximal with respect to inclusion.
@@
An induced subgraph
@@
is called a Dirichlet subgraph if
$\begin{cases}\kappa_{\textnormal{dir}}(y)\coloneqq\kappa_{|Y}(y)+b_{\textnormal{dir}}(y),\\ b_{\textnormal{dir}}(y)\coloneqq\sum_{z\in\accentset{\hbox{\textbullet}}{\partial}Y}w(y,z)=\sum_{z\not\in Y}w(y,z).\end{cases}$ (2.3)
@@ 2.1 Model Problem 2
Let $Y\subsetneq X$ be a nonempty proper subset of the discrete node set $X$ and let $\accentset{\hbox{\textbullet}}{\partial}Y$ be the external boundary of $Y$. In this section we show that the following initial boundary value problem on $Y$
@@
can be reformulated as the (Model Problem 1) over a suitable subgraph.
@@
defines the action of the graph Laplacian associated with the graph $G^{\prime}:=\left(Y,w^{\prime},\kappa_{\operatorname{dir}},\mu^{\prime}\right)\subset G$, where $\kappa_{\operatorname{dir}}$ is as in (2.3), $w^{\prime}=w_{|Y}$ and $\mu^{\prime}=\mu_{|Y}$, we conclude that $u$ solves (Model Problem 1) over $G^{\prime}$ with $h$ as in (2.7).
Let us notice that the Dirichlet BCs in (Model Problem 2) have been absorbed by the forcing term $h$ and the potential term $\kappa_{\operatorname{dir}}$, and that $h\equiv 0$ if $g\equiv 0$.
@@ 2.2 Accretivity
The statement follows from [7, Proposition 2.5] for $p>1$ and from [7, Proposition B.2] with $v=0$ and $\Phi={\rm{id}}$, for $p=1$, noting that in both cases the argument depends on a Green’s formula which is valid for finitely supported function. ∎
@@
This is well known, see, e.g., [7, Proposition 2.5]. We sketch the proof for the convenience of the reader. Accretivity is an immediate consequence of the linearity of $\Delta$ together with (2.9) and Lemma 2.6. Since accretivity implies the injectivity of ${\rm{id}}+\lambda\Delta$ then it is also surjective since it is linear and $C(X)$ is finite dimensional. ∎
@@ 3.1 Definition of solutions
where $h\in L^{1}_{\rm{loc}}([0,T];\ell^{p}\left(X,\mu\right))$ with $1\leq p<\infty$ and $u_{0}\in{\operatorname{dom}_{p}(\mathcal{A})}$.
@@
Mild solutions as defined above are also known in the literature as $C^{0}$ solutions, see [34, Chapter IV.8]. Every strong solution is a mild solution, see e.g., [3, Chapter 4], the converse is true under suitable assumptions, see Theorem 3.7 below.
It is noteworthy to recall that in the literature another notion of mild solution, given in terms of semigroup and inspired by the Duhamel’s principle, is available (see, e.g., [30, Definition 2.3, Chapter 4], [17, Definition 3.1]). We refer to [4, Theorem 5.7] for a proof of the equivalence of the two notions in the linear setting.
@@ 3.2 Existence, uniqueness and regularity of mild solutions
In this section we state the main results about existence, uniqueness and regularity of mild solutions to Problem (3.1) with $\mathcal{A}$ as defined in (2.2) where the operator $F$ is defined as Fu=fuFu=-fu with $f:\mathbb{R}\rightarrow\mathbb{R}$ satisfying one of the following conditions:
@@
$\left\|u(t_{2})-\hat{u}(t_{2})\right\|_{p}\leq\left\|u(t_{1})-\hat{u}(t_{1})\right\|_{p}+\int_{t_{1}}^{t_{2}}\left\|h(s)-\hat{h}(s)\right\|_{p}\,ds,\quad\forall\,0\leq t_{1}<t_{2}\leq T.$ (3.5)
@@ 4 A related time-independent equation
    If (F2) holds with constant $L>0$, then the above statements hold for $\lambda\in(0,1/L)$ with $C=(1-\lambda L)^{-1}$.
@@
As mentioned in the Introduction, the proof of Theorem 4.1 relies on an approximation scheme defined by considering an increasing exhaustion of the graph $X$ by finite sets $\{X_{n}\}_{n}$. The idea is to consider a sequence of operators $\mathcal{A}_{n}$ and data $g_{n}$ defined on finite Dirichlet subgraphs with nodes sets $X_{n}$ and to solve, on each of them, the equation $(\operatorname{id}+\lambda\mathcal{A}_{n})u_{n}=g_{n}$ in $X_{n}$. The solutions $\{u_{n}\}_{n}$ form a sequence that converges, monotonically in case (I) or by dominated convergence in case (II), to a solution of (4.1) which satisfies the required norm estimates. The accretivity (or $\omega$- accretivity) of $\mathcal{A}_{|\Omega_{p}}$ on a dense subset ensures uniqueness. A key role in the proof of the accretivity issue is played by the definition of the set $\Omega_{p}$, which consists of functions in $\text{dom}_{p}(A)$ that can be approximated by sequences supported on the finite sets $X_{n}$, see Definition 4.8 below. Then, the accretivity of $\mathcal{A}$ is recovered by exploiting that of the approximating operators $\mathcal{A}_{n}$. The additional assumptions (IP) or (B) in part (II) are technical and guarantee injectivity of $\text{id}+\lambda\mathcal{A}$ on the full domain, this ensures the convergence of the whole approximating sequence and enables us to replace $\Omega_{p}$ with $\text{dom}_{p}(\mathcal{A})$.
@@ 4.1 Comparison principle
Let $G$ be a connected graph. Suppose that $F$ satisfies (F1) and $\lambda>0$ or that $F$ satisfies (F2) and $0<\lambda<1/L$. Let $u\in\operatorname{dom}\left(\Delta\right)$ and assume that one of the following cases holds:
@@
Define $\psi(t):=t-\lambda f(t)$. If $F$ satisfies (F1) then $\psi$ is strictly monotone increasing and surjective for all $\lambda>0$, while if $F$ satisfies (F2) this holds for all $0<\lambda<1/L$. Indeed, if $0<\lambda<1/L$, then for any $t>s$ one has $\psi(t)-\psi(s)=t-s-\lambda[f(t)-f(s)]\geq(1-\lambda L)(t-s)>0$. Hence, $\psi$ is strictly increasing. Furthermore, we have $(1-\lambda L)t\leq\psi(t)=t-\lambda f(t)\leq(1+\lambda L)t$ for all $t\in\mathbb{R}$, therefore $\lim_{\pm\infty}\psi(t)=\pm\infty$ for all $0<\lambda<1/L$. Hence, $\psi$ is surjective in $\mathbb{R}$ as well.
Let $u\in\operatorname{dom}\left(\Delta\right)$ be such that $\left({\operatorname{id}}+\lambda\mathcal{A}\right)u=\left(\psi+\lambda\Delta\right)u\geq 0$. If $u\geq 0$, then there is nothing to prove. Hence, we assume that there exists $x_{0}\in X$ such that $u(x_{0})<0$. We will show that this leads to a contradiction in the two cases.
@@
Let $\psi$ be defined as in the proof of Theorem 4.3; since it is strictly increasing and $\psi(0)=0$, there exists a positive function $\sigma\in C(X)$ such that
@@ 4.2 Preliminary results on finite graphs
We first prove $(i)$. According to (2.9), with
@@
This concludes the proof of $(i)$. Now $(ii)$ is a consequence of Proposition 2.7 and [4, Example 2.2]. ∎
@@
where $\psi(t):=t-\lambda f(t)$. In the proof of Theorem 4.3 we showed that if $F$ satisfies (F1) then $\psi$ is strictly monotone increasing and surjective for all $\lambda>0$, while if $F$ satisfies (F2) this holds for all $0<\lambda<1/L$.
Then, since $\lambda\Delta$ is a diagonally dominant matrix, all assumptions of [37, Theorem 1] are satisfied and we conclude that, for every $g$, there exists a solution $u$ of (4.7) and, in turn, of (4.1).
@@ 4.3 The set $\Omega_{p}$
As in the proof of Lemma 4.6, we set $\psi(t)=t-\lambda f(t)$. If $F$ satisfies (F1), then $\psi$ is strictly monotone increasing and surjective for every $\lambda>0$, while if $F$ satisfies (F2) the same holds for all $0<\lambda<1/L$. For these values of $\lambda$, from the proof of Lemma 4.6 we know that for every $n\in\mathbb{N}$ there exist $\hat{u}_{n}\in C(X_{n})_{n}$ such that $\hat{u}_{n}\geq 0$ and
@@
If condition (F2) holds, then $F$ maps $\ell^{p}(X,\mu)$ into itself, so that $\operatorname{dom}_{p}(\mathcal{A})=\operatorname{dom}_{p}(\Delta)$. Since $Lt+F(t)$ is monotonically increasing, the above computation shows that
@@ 5 Proofs of the main results
Proposition 5.1 applies in the case when $F$ satisfies condition (F2) and $\sup_{x\in X}\frac{\sum_{x\in X}w(x,y)+\kappa(x)}{\mu(x)}<\infty$. Indeed, it is known [20, Theorem 2.15] that if $\sup_{x\in X}\frac{\sum_{x\in X}w(x,y)+\kappa(x)}{\mu(x)}<\infty$, then $\Delta$ is continuous on $\ell^{p}(X,\mu)$. Moreover, by condition (F2) if $u\in\ell^{p}(X,\mu)$, then $\sum_{x\in X}|F(u(x))|^{p}\mu(x)\leq L^{p}\sum_{x\in X}|u(x)|^{p}\mu(x)$ and if $\{u_{n}\}$ is convergent to $u$ in $\ell^{p}(X,\mu)$, then
@@ 6.2 The linear case f(u)=uf(u)=-u
In this case ${\mathcal{A}}=\Delta+\operatorname{id}$ and Problem (3.1) reads
@@
Still exploiting the accretivity and density results proved in Theorem 4.1 we obtain the following.
Proposition 6.2.
Let $1\leq p<\infty$. Furthermore, if $p>1$ assume that condition ($C_{p}$) holds. If $h\in L^{1}_{\rm{loc}}([0,T];\ell^{p}\left(X,\mu\right))$ and $u_{0}\in\ell^{p}\left(X,\mu\right)$, then Problem (6.1) admits a unique mild solution on $[0,T]$ given by
@@
In particular, it holds
@@
Proof.
By Theorem 4.1, the operator $\mathcal{\mathcal{A}}_{|\Omega_{p}}$ is $m$-accretive and $\Omega_{p}$ is dense in $\operatorname{dom}_{p}(\mathcal{A})$. Furthermore, $\cal{A}$ is linear. Therefore, all assumptions of [4, Theorem 5.7] are satisfied, from which the existence and uniqueness of a mild solution, as given in the statement, follow. Finally, the stated $\ell^{p}$ estimate is a consequence of the $\ell^{p}(X,\mu)$-contractivity of the semigroup associated with $\Delta$, see [20, Theorem 2.9]. ∎

### v3

@@ 1 Introduction
In particular, global existence and blow-up of solutions for parabolic equations has been studied by several authors in [17, 35, 41, 40, 30, 24, 25, 37, 18].
@@
By contrast, Condition (F2) is characteristic of saturating nonlinear forcing terms, which frequently arise in biological models (see, e.g., [8]). For a detailed treatment of this case, see also [43, Chapter 31.6].
@@ 2 Preliminaries and notations
If the cardinality of the node set is finite, i.e., $|X|<\infty$, then $G$ is called a finite graph, otherwise, $G$ is called an infinite graph. We say that two nodes $x$ and $y$ in $X$ are neighbors or connected by an edge, and we write $x\sim y$, if and only if $w(x,y)>0$. We will also say, as in standard graph theory, that $\{x,y\}$ is an edge of the graph. The non-zero values $w(x,y)$ of the edge-weight function $w$ are called weights associated with the edge $\{x,y\}$. On the other hand, if $w(x,y)=0$, then we will write $x\nsim y$ meaning that $x$ and $y$ are not connected by an edge. A walk is a (possibly infinite) sequence of nodes $\{x_{i}\}_{i\geq 0}$ such that $x_{i}\sim x_{i+1}$. A path is a walk with no repeated nodes. A graph is connected if there is a finite walk connecting every pair of nodes, that is, for any pair of nodes $x,y$ there is a finite walk such that $x=x_{0}\sim x_{1}\sim\cdots\sim x_{n}=y$. Moreover, we will say that a subset $Y\subseteq X$ is connected if for every pair of nodes $x,y\in Y$ there exists a finite walk connecting $x$ and $y$ all of whose nodes are in $Y$. A subset $Y\subseteq X$ is a connected component of $X$ if $Y$ is maximal with respect to inclusion.
@@
The presence of $w$ and $\kappa$ gives a one-to-one correspondence between graphs and Markov processes. The edge-weight function $w$ encodes how the process jumps between nodes of $X$ while the killing term $\kappa$ captures the possibility of the process leaving $X$. In other words, the value of $\kappa$ indicates the weight of a connection from that vertex to some additional point outside of $X$. This motivates the name killing term. For a detailed discussion on the meaning of $w$ and $\kappa$ we refer to [21, Part 0].
@@
Given a graph $G=(X,w,\kappa,\mu)$ and $Y\subset X$, we define
[…]
The induced subgraph
@@
is called a Dirichlet subgraph.
@@ 2.1 Model Problem 2
Let $Y\subsetneq X$ be a nonempty proper subset of the discrete node set $X$ and let $\accentset{\hbox{\textbullet}}{\partial}Y$ be the exterior boundary of $Y$. In this section we show that the following initial boundary value problem on $Y$
@@
can be reformulated as (Model Problem 1) by replacing $G$ with a suitable subgraph $G^{\prime}$ and for a suitable choice of the forcing term $h$ absorbing the Dirichlet BCs.
@@
defines the action of the graph Laplacian associated with the graph $G^{\prime}:=\left(Y,w^{\prime},\kappa_{\operatorname{dir}},\mu^{\prime}\right)\subset G$, where $\kappa_{\operatorname{dir}}$ is as in (2.3), $w^{\prime}=w_{|Y}$ and $\mu^{\prime}=\mu_{|Y}$, we conclude that any solution to (Model Problem 2) solves (Model Problem 1) with $G$ replaced by $G^{\prime}$ and $h$ as in (2.7). In particular, $h\equiv 0$ if $g\equiv 0$.
@@ 2.2 Accretivity
The statement follows from [7, Proposition 2.5] for $p>1$ and from [7, Proposition B.2] with $v=0$ and $\Phi={\rm{id}}$, for $p=1$, noting that in both cases the argument depends on a Green’s formula which is valid for finitely supported functions. ∎
@@
This result is well known and follows from the Hille–Yosida Theorem, since $\Delta$ generates a contractive semigroup on $\ell^{p}(X,\mu)$ on finite graphs ([21, Theorem 2.9]). For the reader’s convenience, we briefly sketch the argument. Accretivity is an immediate consequence of the linearity of $\Delta$ together with (2.9) and Lemma 2.6. Since accretivity implies the injectivity of ${\rm{id}}+\lambda\Delta$ then it is also surjective since it is linear and $C(X)$ is finite dimensional. ∎
@@ 3.1 Definition of solutions
where $h\in L^{1}_{\rm{loc}}([0,T];\ell^{p}\left(X,\mu\right))$ with $1\leq p<\infty$ and $u_{0}\in{\ell^{p}(X,\mu)}$.
@@
Mild solutions as defined above are also known in the literature as $C^{0}$ solutions, see [36, Chapter IV.8]. Every strong solution is a mild solution, see e.g., [3, Chapter 4], the converse is true under suitable assumptions, see Theorem 3.7 below.
It is noteworthy to recall that in the literature another notion of mild solution, given in terms of semigroup and inspired by the Duhamel’s principle, is available (see, e.g., [32, Definition 2.3, Chapter 4], [17, Definition 3.1]). We refer to [4, Theorem 5.7] for a proof of the equivalence of the two notions in the linear setting.
@@ 3.2 Existence, uniqueness and regularity of mild solutions
In this section we state the main results about existence, uniqueness and regularity of mild solutions to Problem (3.1) with $\mathcal{A}$ and $F$ defined as in (2.2), with $f:\mathbb{R}\rightarrow\mathbb{R}$ satisfying one of the following conditions:
@@
$\left\|u(t_{2})-\hat{u}(t_{2})\right\|_{p}\leq e^{\omega(t_{2}-t_{1})}\left\|u(t_{1})-\hat{u}(t_{1})\right\|_{p}+\int_{t_{1}}^{t_{2}}e^{\omega(t_{2}-s)}\left\|h(s)-\hat{h}(s)\right\|_{p}\,ds,\quad\forall\,0\leq t_{1}<t_{2}\leq T,$ (3.5)
with $\omega=0$, resp. $\omega=L$, if condition (F1), resp. (F2), hold.
@@ 4 A related time-independent equation
    If (F2) holds with constant $L>0$, then for every $\lambda\in(0,1/L)$ and $g\in\ell^{p,\pm}(X,\mu)$, then there exists a unique solution to equation (4.1) satisfying $u\in\ell^{p,\pm}(X,\mu)\cap\Omega_{p}$ and the estimate
    $\|u\|_{p}\leq C\|g\|_{p},$ (4.3)
    with $C=(1-\lambda L)^{-1}$.
@@
As mentioned in the Introduction, the proof of Theorem 4.1 relies on an approximation scheme defined by considering an increasing exhaustion of the graph $X$ by finite sets $\{X_{n}\}_{n}$. The idea is to consider a sequence of operators $\mathcal{A}_{n}$ and data $g_{n}$ defined on finite Dirichlet subgraphs with nodes sets $X_{n}$ and to solve, on each of them, the equation $(\operatorname{id}+\lambda\mathcal{A}_{n})u_{n}=g_{n}$ in $X_{n}$. The solutions $\{u_{n}\}_{n}$ form a sequence that converges, monotonically in case (I) or by dominated convergence in case (II), to a solution of (4.1) which satisfies the required norm estimates. The accretivity (or $\omega$- accretivity) of $\mathcal{A}_{|\Omega_{p}}$ on a dense subset ensures uniqueness. A key role in the proof of the accretivity issue is played by the definition of the set $\Omega_{p}$, which consists of functions in $\text{dom}_{p}(A)$ that can be approximated by sequences supported on the finite sets $X_{n}$, see Definition 4.8 below. Then, the accretivity of $\mathcal{A}$ is recovered by exploiting that of the approximating operators $\mathcal{A}_{n}$. The additional assumptions (IP) or (B) in part (II) are technical and guarantee injectivity of $\text{id}+\lambda\mathcal{A}$ on the full domain, this ensures the convergence of the whole approximating sequence and enables us to replace $\Omega_{p}$ with $\text{dom}_{p}(\mathcal{A})$. It is worth noting that the exhaustion method adopted here seems the most natural for preserving the actual geometry of the problem. Indeed, the approximation is spatial and dictated by the combinatorial structure of the graph. In particular, this differs from the classical Galerkin method, which relies on arbitrary finite‑dimensional subspaces (typically spanned by a chosen basis) that do not necessarily reflect the geometry of the graph.
@@ 4.1 Comparison principle
Let $G$ be a connected graph. Let $F$ be defined as in (2.2), and suppose that $f$ satisfies (F1) and $\lambda>0$ or $f$ satisfies (F2) and $0<\lambda<1/L$. Let $u\in\operatorname{dom}\left(\Delta\right)$ and assume that one of the following cases holds:
@@
Define $\psi(s):=s-\lambda f(s)$, $s\in\mathbb{R}$. If $f$ satisfies (F1) then $\psi$ is strictly monotone increasing and surjective for all $\lambda>0$, while if $f$ satisfies (F2) this holds for all $0<\lambda<1/L$. Indeed, if $0<\lambda<1/L$, then for any $s>s^{\prime}$ one has $\psi(s)-\psi(s^{\prime})=s-s^{\prime}-\lambda[f(s)-f(s^{\prime})]\geq(1-\lambda L)(s-s^{\prime})>0$. Hence, $\psi$ is strictly increasing. Furthermore, we have $|\psi(s)|=|s-\lambda f(s)|\geq(1-\lambda L)|s|$ for all $s\in\mathbb{R}$, therefore $\lim_{\pm\infty}\psi(s)=\pm\infty$ for all $0<\lambda<1/L$. Hence, $\psi$ is surjective in $\mathbb{R}$ as well. Let $u\in\operatorname{dom}\left(\Delta\right)$ be such that $\left({\operatorname{id}}+\lambda\mathcal{A}\right)u=\left(\psi+\lambda\Delta\right)u\geq 0$. If $u\geq 0$, then there is nothing to prove. Hence, we assume that there exists $x_{0}\in X$ such that $u(x_{0})<0$. We will show that this leads to a contradiction in the two cases.
@@
Let $\psi$ be defined as in the proof of Theorem 4.3; since it is strictly increasing and $\psi(0)=0$, $\psi(a)-\psi(b)$ and $\psi(a-b)$ have the same sign for all $a,b\in\mathbb{R}$. Hence, there exists a positive function $\sigma\in C(X)$ such that
@@ 4.2 Preliminary results on finite graphs
Both statements follow from general theory. Statement $(ii)$ is proved in [4, Example 2.2]. As for $(i)$, it can be obtained by combining [4, Proposition 2.20] and [4, Proposition 2.23]. For the reader’s convenience, we nevertheless include below a direct elementary proof well suited to our setting.
According to (2.9), with
@@
This concludes the proof of $(i)$.
@@
where $\psi(s):=s-\lambda f(s)$. In the proof of Theorem 4.3 we showed that if $f$ satisfies (F1) then $\psi$ is strictly monotone increasing and surjective for all $\lambda>0$, while if $f$ satisfies (F2) this holds for all $0<\lambda<1/L$.
Then, since $\lambda\Delta$ is a diagonally dominant matrix, all assumptions of [39, Theorem 1] are satisfied and we conclude that, for every $g$, there exists a solution $u$ of (4.8) and, in turn, of (4.1).
@@ 4.3 The set $\Omega_{p}$
As in the proof of Lemma 4.6, we set $\psi(s)=s-\lambda f(s)$. If $f$ satisfies (F1), then $\psi$ is strictly monotone increasing and surjective for every $\lambda>0$, while if $f$ satisfies (F2) the same holds for all $0<\lambda<1/L$. For these values of $\lambda$, from the proof of Lemma 4.6 we know that for every $n\in\mathbb{N}$ there exist $\hat{u}_{n}\in C(X_{n})_{n}$ such that $\hat{u}_{n}\geq 0$ and
@@
If condition (F2) holds, then $F$ maps $\ell^{p}(X,\mu)$ into itself, so that $\operatorname{dom}_{p}(\mathcal{A})=\operatorname{dom}_{p}(\Delta)$. Since $s\mapsto Ls-f(s)$ is monotonically increasing, the above computation shows that
@@ 5 Proofs of the main results
Proposition 5.1 applies in the case when $f$ satisfies condition (F2) and $\sup_{x\in X}\frac{\sum_{y\in X}w(x,y)+\kappa(x)}{\mu(x)}<\infty$. Indeed, it is known [21, Theorem 2.15] that if $\sup_{x\in X}\frac{\sum_{y\in X}w(x,y)+\kappa(x)}{\mu(x)}<\infty$, then $\Delta$ is continuous on $\ell^{p}(X,\mu)$. Moreover, by condition (F2) if $u\in\ell^{p}(X,\mu)$, then $\sum_{x\in X}|F(u(x))|^{p}\mu(x)\leq L^{p}\sum_{x\in X}|u(x)|^{p}\mu(x)$ and if $\{u_{n}\}$ is convergent to $u$ in $\ell^{p}(X,\mu)$, then
@@ 6.2 The linear case f(u)=uf(u)=-u
In this case ${\mathcal{A}}=\Delta+\operatorname{id}$ is linear and Problem (3.1) reads
@@
Let $1\leq p<\infty$. Furthermore, if $p>1$ assume that condition ($C_{p}$) holds. By Theorem 4.1, the operator $\mathcal{\mathcal{A}}_{|\Omega_{p}}$ is $m$-accretive and $\Omega_{p}$ is dense in $\operatorname{dom}_{p}(\mathcal{A})$.
Therefore, all assumptions of [4, Theorem 5.7] are satisfied, from which, for all $h\in L^{1}_{\rm{loc}}([0,T];\ell^{p}\left(X,\mu\right))$ and $u_{0}\in\ell^{p}\left(X,\mu\right)$, Problem (6.1) admits a unique mild solution on $[0,T]$ given by
@@
Finally, as a consequence of the $\ell^{p}(X,\mu)$-contractivity of the semigroup associated with $\Delta$ (see [21, Theorem 2.9]) $u$ satisfies